OpenStudy (kaylak12345):

in which quadrant does this solution fall y=x-1 and 2x-y=2 @SolomonZelman

OpenStudy (kaylak12345):

sorry wrong one hold up

OpenStudy (solomonzelman):



OpenStudy (solomonzelman):

sure...

OpenStudy (kaylak12345):

y=x-1 and -3x-5

OpenStudy (solomonzelman):

\(\Large\color{black}{y=x-1 }\) and \(\Large\color{black}{y=-3x-5 }\) right?

OpenStudy (kaylak12345):

yes quadrant 3 right

OpenStudy (solomonzelman):

a solution is a point in which the 2 equation intersect. In other words, a value for x and y that makes both of the equations be true.

OpenStudy (solomonzelman):

So, lets solve the equations. I would do substitution.

OpenStudy (kaylak12345):

ok

OpenStudy (solomonzelman):

You are given that y is equal to \(\Large\color{black}{x-1 }\) and at the same time the y is also equal to \(\Large\color{black}{-3x-5 }\). this way, we can say that \(\Large\color{black}{x-1=-3x-5 }\) . just like saying, if a=b, and a=c, THEN b=c.

OpenStudy (kaylak12345):

so it is quadrant 3 because thats where the points intersect

OpenStudy (solomonzelman):

no.

OpenStudy (solomonzelman):

can yo solve for x, when we have \(\Large\color{black}{x-1=-3x-5 }\) ? and do you know how we arrive at this?

OpenStudy (kaylak12345):

it isnt really thats what I am getting according to my graphing calculator

OpenStudy (solomonzelman):

I don't know what you are getting according to your graphing calculator, but you are supposed to solve the equation and prove it without using a graphing calculator. and this is where I am trying to get.

OpenStudy (solomonzelman):

so do you understand everything so far, up to \(\Large\color{black}{x-1=-3x-5 }\) ?

OpenStudy (kaylak12345):

ok and yes

OpenStudy (kaylak12345):

without using a graphing calculator I am still getting quadrent 3

OpenStudy (solomonzelman):

now, \(\Large\color{black}{x-1=-3x-5 }\) \(\Large\color{black}{x-1\color{blue}{ +1 }=-3x-5 \color{blue}{ +1 }}\) \(\Large\color{black}{x=-3x-4}\) \(\Large\color{black}{x\color{blue}{ +3x }=-3x-4\color{blue}{ +3x }}\) \(\Large\color{black}{4x=-4}\)

OpenStudy (solomonzelman):

Find x.

OpenStudy (solomonzelman):

After finding x, find the y....

OpenStudy (kaylak12345):

I have the answrt and madre a 100

OpenStudy (solomonzelman):

cool....

OpenStudy (kaylak12345):

yep